# Kernel Search approach to solve the Minimum Spanning Tree Problem with conflicting edge pairs00footnotetext: Email addresses:

### v1

@@ Abstract
The Minimum Spanning Tree Problem with Conflicts consists in finding the minimum conflict-free spanning tree of a graph, i.e., the spanning tree of minimum cost, including no pairs of edges that are in conflict. In this paper, we solve this problem using a tailored Kernel Search heuristic method, which consists in solving iteratively improved restrictions of the problem. The main novelty of the approach consists in using an independent set of the conflict graph within the algorithm. We test our approach on the benchmark instances and we compare our results with the ones obtained by other heuristics available in the literature.
Keywords: Minimum Spanning Tree, Conflicting Edges, Kernel Search, Cybersecurity, Disjunctive Constraints, Matheuristic
@@ 1 Introduction
In this paper, we propose a new solution approach to the MSTC based on the so-called Kernel Search (KS) method [1]. It falls within the category of general-purpose methods for solving Mixed-Integer Linear Programming (MILP) problems. Initially employed in addressing challenges like the multidimensional knapsack problem [1] and a portfolio selection problem [2], it has been used also to solve, among others, facility location problems with or without conflicts [25, 19, 21], as well as the multivehicle inventory routing problem [3]. The main purpose of this algorithm is to obtain a feasible solution, hopefully of good quality, from a small set of promising variables of the MILP. This subset of variables constitutes the so-called kernel set that is initially built using the information provided by solving the Linear Programming (LP) relaxation of the original MILP. Throughout the algorithm, new promising variables to add to the kernel set may be found by solving MILP subproblems of increasing size.
An important novelty in our approach relies on the use of independent sets of proper subgraphs of the conflict graph, within the KS algorithm. This idea is derived from the observation that the set of edges of any feasible solution of MSTC corresponds to an independent set in the conflict graph, and we believe that it can be applied to other optimization problems with conflicts, whenever the conflict graph is used.
The rest of the paper is organized as follows. In Section 2, we present the Integer Linear Programming (ILP) formulation of MSTC we consider. In Section 3, the tailored kernel search approach we use to solve this formulation is proposed and discussed. In Section 4, we test the KS approach on the benchmark instances and compare our results with the ones obtained by other heuristics available in the literature. Finally, Section 5 concludes the paper.
@@ 2 ILP Formulation of MSTC and notation
In the following, we denote by LP-MSTC the LP relaxation of model (2), obtained by relaxing constraints (1e). Given a subset of edges $F\subseteq E$, we further denote by MSTC$(F)$ the ILP restriction of problem (2) obtained by restricting the set of variables to the edges in $F$, i.e., adding constraints
@@
to formulation (2). As before, LP-MSTC$(F)$ denotes the LP relaxation of MSTC$(F)$.
@@
To conclude this section, we recall some important definitions. The set $N(u)$ of the neighbors of node $u\in V$ is the set of all the nodes $i\in V\setminus\{u\}$ such that it exists an edge $\{u,v\}\in E$. An independent set is a set of nodes in a graph, no two of which are adjacent. Given the graph $\mathcal{G}=(V,E)$, its conflict graph $\hat{\mathcal{G}}=(\hat{V},\hat{E})$ is a graph which has a node for each edge in $\mathcal{G}$, and where we represent each conflict constraint by an edge connecting the corresponding vertices in $\hat{\mathcal{G}}$. Note that each conflict-free spanning tree in $\mathcal{G}$ is a subset of $\hat{V}$ which corresponds to an independent set of $\hat{\mathcal{G}}$.
3 Kernel Search approach
In this section, we describe the KS framework we use to solve MSTC. Given the mathematical formulation (2) of MSTC, the starting point of the KS algorithm consists in identifying a subset of promising variables of this model, i.e., variables that have a high probability of being non-zero in the optimal solution of the problem. These variables are used to form a subset called kernel set, while the remaining ones are partitioned into several subsets called buckets. We define $\Lambda^{0}$ as the set of indices of the variables in the initial kernel set, and $B_{i}$ as the set of indices of the variables in the $i$-th bucket, for $i=1,\dots,b$.
The partitioning of the indices of the variables into the kernel set and the buckets is carried out by KS according to the optimal solution of LP-MSTC, an MST solution (not necessarily conflict-free) returned by the GRASP algorithm proposed in [14], as well as an independent set of a proper subgraph of the conflict graph $\hat{\mathcal{G}}$. After the construction of the kernel set and the buckets, KS seeks an initial feasible solution by solving the MSTC restricted to the variables having indices inside the kernel set only, i.e. solving MSTC($\Lambda^{0}$). For this reason, the initial kernel size $|\Lambda^{0}|$ should be, at the same time, large enough to contain as many variables as possible among the ones that are likely to be part of an optimal solution, and not too large to ensure that MSTC($\Lambda^{0}$) can be efficiently solved to optimality.
The kernel set is then updated throughout the KS algorithm ($\Lambda^{k}$ is the kernel set at the end of iteration $k$), by taking into account the buckets. At each iteration, indeed, the MILP formulation, restricted to the current kernel set and a bucket from the sequence, is solved, trying to obtain solutions of increasing quality and thus new variables to insert into the kernel set.
@@ 3 Kernel Search approach
  • Extension phase: At each iteration $k$, new variables are added to the kernel set iterating (eventually more than once) over the buckets, and the MSTC restricted on the so-obtained enlarged kernel is solved. If this restriction is feasible, its optimal value is denoted by $w^{k},$ and eventually used to update the value of $w^{*}$, and its optimal solution is denoted by $x^{k},$ and eventually used to update the kernel set and the buckets.
Let us describe in more detail how these two phases work.
@@ 3.1 Initialization phase
The construction of starting kernel set $\Lambda^{0}$ and of set $\mathcal{B}$ is carried out by Algorithm 2. The algorithm takes as input two positive parameters $K$ and $d$. First of all, the optimal solution of LP-MSTC is computed (line 1). The variables that have a positive value in the solution of LP-MSTC are then sorted according to their values, in non-increasing order (line 2). Being $N$ the set of edge indices associated with the positive variables ordered in this way, we set $K$ in such a way that it is not greater than the cardinality of $N$ (line 3). We define the subset of the first $K$ elements of $N$ as $N_{K}$ and the set of the pairs of edges in $N_{K}$ which are in conflict as $C_{K}$, i.e., $C_{K}=\{(i,j)~{}:~{}i,j\in N_{K},\text{ and }(e_{i},e_{j})\in C\}$ (line 4). The graph $\hat{\mathcal{G}}_{K}=(N_{K},C_{K})$ (subgraph of the conflict graph $\hat{\mathcal{G}}$) is taken into account. In line 5, a feasible solution $I$ of the maximum independent set problem on this graph $\hat{\mathcal{G}}_{K}$, with the guarantee of not generating cycles in the starting graph $\mathcal{G}$, is computed through a greedy procedure presented in Algorithm 1, by setting $W=N_{K}$ and $\mathcal{H}=\hat{\mathcal{G}}_{K}$.
We use the maximum independent set problem to find the most promising variables by reason of the fact that a spanning tree $(V,E^{\prime})$ is a feasible solution for the MSTC problem iff $E^{\prime}$ is an independent set of the conflict graph. Notice that this is a necessary but not sufficient condition to obtain a feasible solution for the MSTC. Since Algorithm 1 selects edges that are not in conflict with each other and that do not generate cycles in the starting graph $\mathcal{G}$, then these edges induce either a feasible solution of the MSTC or a set of conflict-free subtrees of $\mathcal{G}$.
This means that, by selecting the variables according to the solution of Algorithm 1, we increase the probability of including edges in the kernel set that can generate feasible solutions for the MSTC. It is worth noting that, actually, the idea of using the maximum independent set can be extended to any optimization problem with conflicts because its feasible solutions must be independent sets of the conflict graph.
Input: Graph $\mathcal{H}$, and node set $W$. Output: Independent set $I$ of $\mathcal{H}$. 1 Set $I=\emptyset,\bar{V}=\emptyset,\bar{E}=\emptyset$. 2 while $W\neq\emptyset$ do 3       Select a node $i\in W$ with minimum degree in $\mathcal{H}[W]$, and let $e_{i}=\{u,v\}\in E$ be the edge of $\mathcal{G}$ associated to $i$. 4       Set $\bar{V}=\bar{V}\cup\{u,v\}$ and $\bar{E}=\bar{E}\cup\{e_{i}\}.$ 5       if  $\mathcal{G}[\bar{E}]$ does not contain cycles then 6            Set $I=I\cup\{i\}.$ 7             Being $W_{i}$ the set of neighbors of $i$ in $\mathcal{H}$, including $i$, set $W=W\setminus W_{i}.$ 8      else 9             Set $W=W\setminus\{i\}.$ 10       end if 11       12 end while return $I$ Algorithm 1 Algorithm to compute an independent set of the graph $\mathcal{H}$
Once the set $I$ is computed through Algorithm 1, the GRASP algorithm proposed in [14] is executed (line 5 as well). This multistart metaheuristic consists of two phases: (i) a constructive phase, which creates an initial solution using the adaptive memory (i.e., taking into account a measure of how frequently an edge occurs in the best solutions generated in previous iterations); (2) a local search procedure, moving from the current solution to neighbor ones, trying to find better solutions. The output of this algorithm is a tree $(V,\bar{E})$, not necessarily without conflicts.

### v2

# Advanced Kernel Search approach for the MST Problem with conflicts involving affinity detection and initial solution construction00footnotetext: *Corresponding author Email addresses:

@@ Abstract
The Minimum Spanning Tree Problem with Conflicts consists in finding the minimum conflict-free spanning tree of a graph, i.e., the spanning tree of minimum cost, including no pairs of edges that are in conflict. In this paper, we solve this problem using an enhanced Kernel Search method, which iteratively solves refined problem restrictions. Our approach addresses two central open questions in the kernel search literature: (1) how to determine the affinity between variables to ensure that the restricted problem contains variables that are as compatible as possible, meaning they are more likely to appear together in a feasible solution, and (2) how to construct an initial feasible solution quickly. To this end, we integrate the computation of independent sets from the conflict graph within the algorithm to detect affinities and effectively manage conflicts. Furthermore, we heuristically construct an initial starting point, significantly accelerating the computational process. Although our methodology is designed for MSTC, its principles could be extended to other combinatorial optimization problems with conflicts. Experimental results on benchmark instances demonstrate the efficiency and competitiveness of our approach compared to existing methods in the literature, achieving 17 new best-known values.
@@ 1 Introduction
In this paper, we propose a new solution approach to the MSTC based on the so-called Kernel Search (KS) method [1]. It falls within the category of general-purpose methods for solving Mixed-Integer Linear Programming (MILP) problems. Initially employed in addressing challenges like the multidimensional knapsack problem [1] and a portfolio selection problem [2], it has also been used to solve, among others, facility location problems with or without conflicts [28, 20, 23], as well as the multivehicle inventory routing problem [3]. The main purpose of this algorithm is to obtain a feasible solution, hopefully of good quality, from a small set of promising variables of the MILP. The algorithm operates in two phases: in the initialization phase, the so-called kernel set is initially built using the information provided by solving the Linear Programming (LP) relaxation of the original MILP. In the subsequent improvement phase, new promising variables to add to the kernel set are iteratively identified by solving MILP subproblems of increasing size.
One of the novelties in our approach is the integration of independent sets (sets of vertices, no two of which are adjacent) from proper subgraphs of the conflict graph within the KS algorithm. This is motivated by the observation that the set of edges in any feasible solution for the MSTC corresponds to an independent set in the conflict graph. By leveraging this in both the initialization and improvement phases, we ensure that the variables included in the reduced problem exhibit high compatibility (or affinity), thereby increasing the likelihood of forming a feasible solution. We believe this concept has the potential to be extended to other optimization problems involving conflicts where the conflict graph is utilized. Another distinguishing aspect of our methodology is the iterative nature of the KS algorithm we devise, in which buckets are not only revisited but also merged based on independent sets identified in the conflict graph. This approach enhances the algorithm ability to manage conflicts effectively and refine the solution progressively. Finally, we introduce a heuristic procedure for constructing a starting point still employing independent sets of the conflict graph. This addresses one of the main challenges of the KS approach: the difficulty of finding a feasible starting solution, particularly for complex problems [22]. These starting solutions are particularly useful for building the initial kernel set for problems like MSTC in which the optimal solution of the continuous relaxation does not provide enough information to identify promising variables. The proposed KS algorithm is tested on benchmark instances, achieving 17 new best-known values, demonstrating its effectiveness in advancing the state of the art.
The rest of the paper is organized as follows. In Section 2, we present the Integer Linear Programming (ILP) formulation of MSTC we consider. In Section 3, we outline the KS framework. In Section 4, we present a heuristic finding an independent set of a given conflict graph (Section 4.1) and another finding a promising set of edges in the graph (Section 4.2). In Section 5, we describe the two phases of the proposed KS approach in detail. A preprocessing procedure as well as some valid inequalities for the MSTC are described, which are used to strengthen the LP relaxation of the problem. In Section 6, we test the KS approach on the benchmark instances and compare our results with the ones obtained by other heuristics available in the literature. Finally, Section 7 concludes the paper.
@@ 2 ILP Formulation of MSTC and notation
In the following, we denote by LP-MSTC the LP relaxation of model (1), obtained by relaxing constraints (1e). In addition, we define as $\overline{\text{LP-MSTC}}$ the model LP-MSTC without the subtour elimination constraints (constraints (1c)). Given a subset of edges $F\subseteq E$, we further denote by MSTC$(F)$ the ILP restriction of problem (1) obtained by restricting the set of variables to the edges in $F$, i.e., adding constraints
@@
to formulation (1). As before, LP-MSTC$(F)$ denotes the LP relaxation of MSTC$(F)$. We further use the notation $conflicts(F)$ to represent the number of conflicts in the set of edges $F$, i.e., the cardinality of set $\{i,j\in F:(e_{i},e_{j})\in C\}$.
@@
To conclude this section, we recall some important definitions. Let $N(u)=\{v\in V:\{u,v\}\in E\}$ be the set of neighbors of node $u$ that is the set of nodes adjacent to $u$ in $\mathcal{G}$. An independent set is a set of nodes in a graph, no two of which are adjacent. Given the graph $\mathcal{G}=(V,E)$, its conflict graph $\hat{\mathcal{G}}=(\hat{V},\hat{E})$ is a graph which has a node for each edge in $\mathcal{G}$, and where we represent each conflict constraint by an edge connecting the corresponding vertices in $\hat{\mathcal{G}}$. Note that each conflict-free spanning tree in $\mathcal{G}$ is a subset of $\hat{V}$ which corresponds to an independent set of $\hat{\mathcal{G}}$.
3 Preliminaries
In this section, we briefly outline the KS framework used to solve the MSTC problem and introduce two algorithms employed within it. Given the mathematical formulation (1) of MSTC, let $\Omega$ be the set of ideal variables, i.e., the variables taking a value equal to one in the optimal solution of MSTC. It is easy to see that the resolution of the subproblem MSTC$(\Omega)$ provides the optimal solution to the MSTC problem: all the $|\Omega|=n-1$ variables must be set to 1. As $\Omega$ is not known, the KS algorithm consists in identifying a set $\Lambda$ of promising variables of this model, i.e., variables that have a high probability of being non-zero in the optimal solution of the problem. This is because, if $\Omega\subseteq\Lambda$, the resolution of MSTC$(\Lambda)$ provides the optimal solution to the MSTC problem. The starting set of promising variables is used to form the kernel set, while the remaining ones are partitioned into several subsets called buckets. We define $\Lambda^{0}$ as the set of variables indices in the initial kernel set, and $B_{k}$ as the set of variables indices in the $k$-th bucket, for $k=1,\dots,b$.
The partitioning of the indices of the variables into the kernel set and the buckets is carried out by KS according to the optimal solution of LP-MSTC, (eventually) a feasible MSTC solution, as well as an independent set of a proper subgraph of the conflict graph $\hat{\mathcal{G}}$. After the construction of the kernel set and the buckets, KS seeks an initial feasible solution by solving the MSTC restricted to the variables having indices inside the kernel set only, i.e., solving MSTC($\Lambda^{0}$). For this reason, the initial kernel size $|\Lambda^{0}|$ should be, at the same time, large enough to contain as many variables as possible among the ones that are likely to be part of an optimal solution, and not too large to ensure that MSTC($\Lambda^{0}$) can be efficiently solved to optimality.
The kernel set is then updated throughout the KS algorithm ($\Lambda^{k}$ is the kernel set at the end of iteration $k$), by taking into account the buckets. At each iteration, indeed, the MILP formulation, restricted to the current kernel set and one or multiple buckets from the sequence, is solved, trying to obtain solutions of increasing quality and thus new variables to insert into the kernel set.
@@ 3 Preliminaries
  • Improvement phase: At each iteration $k$, the MSTC($\Lambda^{k-1}\cup B_{k}$) is solved. If this subproblem is feasible, its optimal solution $x^{k}$ and relative value $w^{k}$ are used to, eventually, update the incumbent solution and its value, respectively. Moreover, all the variables of $B_{k}$, greater than zero in $x^{k}$, are moved into the kernel set, generating the new kernel set $\Lambda^{k}$. The iteration over the buckets may be performed more than once, considering new buckets obtained from the union of the original buckets.
All the parameters that will be involved in the KS algorithm are listed in Table 1.
Parameter Description
$\bar{h}$ maximum number of for-loop iterations of the StartingSolution algorithm
$\bar{t}$ maximum number of while-loop iterations of the StartingSolution algorithm
$b$ number of buckets
$d$ cardinality of each bucket
$K$ cardinality of the set which is used to create the first kernel set
$P$ number of iterations performed over the buckets in the improvement phase
$p\in\{1,P\}$ index of the iteration performed over the buckets in the improvement phase representing the type of bucket
added to the kernel set (outer iterations)
$\Delta$ maximum number of buckets added to the kernel at each iteration, without finding a feasible MSTC solution
$\bar{k}\in\{1,\Delta\}$ index representing the number of buckets already added to the kernel set without finding a feasible MSTC solution
$k\in\{1,b\}$ index representing the bucket added to the kernel set (inner iterations)
$\Lambda^{k}$ kernel set at iteration $k$
$B_{k}$ $k$-th bucket with $k\in\{1,b\}$
$\mathcal{B}$ set of buckets
Table 1: List of the parameters involved in the KS algorithm.
4 Embedded heuristics
The effectiveness of the KS approach depends on the quality of the initial kernel set. This set is usually generated using the information retrieved from the LP relaxation of the model. However, for some problems, this information alone is not sufficient to ensure a high-quality initial kernel set. For this reason, an alternative approach, recently investigated in [5], consists of applying fast heuristic methods to generate feasible solutions to the problem. These solutions provide additional information, which, together with the LP relaxation solutions, can be used to construct the initial kernel set. To this end, in this section, we introduce two heuristics.
The first heuristic identifies independent sets in the conflict graph, ensuring that the selected edges do not induce cycles in the original graph. While this guarantees a conflict-free selection, the resulting subgraph may not be connected.
To improve connectivity, the second heuristic uses the first one but iteratively extends its outputs by incorporating additional edges. It can yield different outcomes: a conflict-free spanning forest, a spanning tree that includes some conflicting edges, or an entirely conflict-free spanning tree. It leverages the first heuristic within its iterations to maintain as many conflict-free edges as possible. Through this approach, the second heuristic effectively attempts to produce a feasible MSTC solution.
4.1 Computing an independent set of the conflict graph
In this section, we describe a greedy algorithm, IndependentSet, whose aim is to compute an independent set $I$ of the conflict graph $\mathcal{\hat{G}}$ that does not create cycles inside the original graph $\mathcal{G}$. The usefulness of this algorithm derives from the following observation.
Observation 1
Any feasible MSTC solution is composed of a set of $n-1$ edges that: (i) does not create cycles in $\mathcal{G}$; (ii) corresponds to an independent set of $\mathcal{\hat{G}}$.
Our idea is to use the IndependentSet algorithm during both phases of KS to find promising variables. It is worth noting that the idea of using the maximum independent set can be extended to any optimization problem with conflicts because its feasible solutions must be independent sets of the conflict graph.
The pseudocode of the IndependentSet algorithm is shown in Algorithm 1. The algorithm takes in input the graphs $\mathcal{G}$ of MSTC and a subgraph $\mathcal{H}$ of the conflict graph $\mathcal{\hat{G}}$. The independent set $I$ is initialized to the empty set on line 1 together with sets $\bar{V}$ and $\bar{E}$. The set of vertices $W$ is initialized with all the vertices of $\mathcal{H}$ (line 1), and the while loop (lines 1-1) is repeated until $W$ is not empty.
Input: Graph $\mathcal{G}=(V,E)$, graph $\mathcal{H}=(V_{H},E_{H})$.
Output: Independent set $I$ of $\mathcal{H}$.
1 $I\leftarrow\emptyset,\bar{V}\leftarrow\emptyset,\bar{E}\leftarrow\emptyset$.
2 $W\leftarrow V_{H}.$
3 while $W\neq\emptyset$ do
4       Select a node $i\in W$ with minimum degree in $\mathcal{H}[W]$, and let $e_{i}=\{u,v\}\in E$ be the edge of $\mathcal{G}$ associated to $i$.
5       if  $\mathcal{G}[\bar{E}\cup\{e_{i}\}]$ does not contain cycles then
6            $I\leftarrow I\cup\{i\}.$
7             $\bar{V}\leftarrow\bar{V}\cup\{u,v\}$, $\bar{E}\leftarrow\bar{E}\cup\{e_{i}\}.$
8             $W_{i}\leftarrow$ set of neighbors of $i$ in $\mathcal{H}$, including $i$.
9            $W\leftarrow W\setminus W_{i}.$
10      else
11             $W\leftarrow W\setminus\{i\}.$
12
return $I$
Algorithm 1 IndependentSet
The algorithm selects a node $e_{i}$ of $W$ having the minimum degree in the induced subgraph $\mathcal{H}[W]$ (line 1). Let us suppose that $e_{i}$ corresponds to the edge $i$ in $\mathcal{G}$ whose extremes are $u$ and $v$. The IndependentSet algorithm checks if the subgraph of $\mathcal{G}$ induced by $\bar{E}\cup\{e_{i}\}$ contains cycles (line 1). If this is the case, the node $i$ is removed from W (line 1). Otherwise, $i$ is added to the independent set $I$ (line 1), and $\bar{E}$ and $\bar{V}$ are updated by adding the edge $e_{i}$ and its extremes $u$ and $v$, respectively (line 1). Finally, $i$ and its neighbors are removed from $W$ (line 1). Once $W$ becomes empty, the algorithm returns the built independent set $I$.
Since the IndependentSet algorithm selects edges that are not in conflict with each other and that do not generate cycles in the starting graph $\mathcal{G}$, then these edges induce either a feasible solution of the MSTC or a set of conflict-free subtrees of $\mathcal{G}$. This means that, by selecting the variables according to the solution of the IndependentSet algorithm, we increase the probability of including edges in the kernel set that can generate feasible MSTC solutions. We further use the IndependentSet algorithm within the improvement phase to decide how to merge the buckets if multiple of them are supposed to be considered at a time.
4.2 Finding a feasible starting point
In this section, we present the StartingSolution heuristic procedure that tries to generate a feasible MSTC solution, leveraging the set of edges returned by the IndependentSet algorithm, which is conflict-free, but not necessarily connected. The StartingSolution pseudo-code is shown in Algorithm 2.
Input: Graph $\mathcal{G}=(V,E)$, conflict graph $\hat{\mathcal{G}}$, initial edge set $S^{0}$, two integer parameters $\bar{h},\bar{t}>1$.
Output: Either a conflict-free spanning tree or an empty set.
1 $S\leftarrow S^{0}$.
2 $\bar{E}\leftarrow\emptyset$.
3 $\bar{w}\leftarrow+\infty$.
4 for $h=1,\dots\bar{h}$  do
5       Launch the Kruskal algorithm on $\mathcal{G}[S]$ obtaining a set of edges $E^{\prime}$.
6       $\hat{E}\leftarrow\{e_{i}\in E^{\prime}:\exists\;e_{j}\in E^{\prime}\;s.t.\;(e_{i},e_{j})\in C\}$.
7       $t\leftarrow 0$
8       while $|\hat{E}|\neq 0$ and $t\leq\bar{t}$  do
9             Compute the independent set $I^{\prime}$ over graph $\hat{\mathcal{G}}[\hat{E}]$ through the IndependentSet algorithm.
10             $E^{\prime}\leftarrow E^{\prime}\setminus(\hat{E}\setminus I^{\prime})$.
11             Compute the independent set $I$ over graph $\hat{\mathcal{G}}$ through the IndependentSet algorithm, imposing that $E^{\prime}\subseteq I$.
12             if $I$ is a spanning tree then
13                  $\hat{E}\leftarrow\emptyset$ and $E^{\prime}\leftarrow I$
14                   break
15            else
16                  Assign a random weight ${w^{\prime}}_{e}$ to each edge $E$.
17                   Connect $I$ with the edges in set $\bar{I}\subseteq E\setminus I$ such that $\sum_{e\in\bar{I}}{w^{\prime}}_{e}$ is minimized, obtaining the tree $T$.
18                   $E^{\prime}\leftarrow T.$
19                   $\hat{E}=\{e_{i}\in T:\exists\;e_{j}\in T\;s.t.\;(e_{i},e_{j})\in C\}$.
20            $t\leftarrow t+1$
21
22      $S\leftarrow S\cup E^{\prime}.$
23       if $\hat{E}=\emptyset$ and $\sum\limits_{e\in E^{\prime}}w_{e}<\bar{w}$ then
24            $\bar{E}\leftarrow E^{\prime}$
25             $\bar{w}\leftarrow\sum\limits_{e\in E^{\prime}}w_{e}$
26             if $t=0$ then break;
27
return $S,\bar{E}$
Algorithm 2 StartingSolution
The procedure makes use of Kruskal’s algorithm [27]. The process starts with a set of edges $S=S^{0}$. A for-loop of maximum $\bar{h}$ iterations begins, and Kruskal’s algorithm is applied to the induced subgraph $\mathcal{G}[S]$ (line 2-2). If the resulting forest $E^{\prime}$ (or tree if $\mathcal{G}[S]$ is connected) contains conflicting edges, a while-loop of at most $\bar{t}$ iterations is initiated (line 2). Let $\hat{E}$ be the edges of $E^{\prime}$ that are in conflict with each other. Within each iteration of this while-loop, an independent set $I^{\prime}$ is computed over the conflict graph obtained from the conflicting edges in $E^{\prime}$, using the IndependentSet algorithm (line 2). The edges in $\hat{E}$ that are not part of $I^{\prime}$ are removed from the forest $E^{\prime}$ (line 2). A new independent set $I$, containing $E^{\prime}$, is computed by the IndependentSet algorithm on the whole conflict graph $\hat{G}$ (line 2). If $I$ forms a spanning tree, a feasible solution to the MSTC problem is obtained, and the while-loop terminates (line 2). Otherwise, additional edges are added to complete the tree through a graph augmentation process (lines 22). This graph augmentation procedure finds a minimum spanning tree $T$, containing $I$, according to randomly assigned weights. If any conflicting edge pair is detected in $T$ (line 2), and the maximum number of iterations for the while-loop has not been reached, a new iteration begins. Otherwise, the conflict-free tree is added to the set $S$ (line 2) and eventually used to update the incumbent solution (lines 22). With this enlarged set $S$, the for-loop then proceeds to the next iteration, unless either the maximum number of iterations has been reached or $t=0$ (line 2). This last condition holds when $\hat{E}=\emptyset$ on line 2 and, thus, the while-loop of line 2 is not performed. Consequently, $E^{\prime}\subseteq S$ and therefore $S$ (line 2) could not change for the remaining iterations of the for-loop. In the end, the heuristic will return as $\bar{E}$ one of the following outcomes: either an empty set or a conflict-free spanning tree.
As the heuristic will be used in the initialization phase of the KS algorithm to create the initial kernel set $\Lambda^{0},$ and the returned set of edges may be also empty, the heuristic further returns the set $S$ of analyzed edges.
5 Kernel Search Approach
We are now ready to describe in detail our KS approach. In Section 5.1, we focus on the initialization phase, which aims at defining the starting kernel set and the starting set of buckets. In Section 5.2, we describe the improvement phase of the algorithm, in which better and better restrictions of the MSTC problem are solved. In the rest of the paper, we will refer to the entire KS algorithm, including both the initialization and extension phases, as KS.
@@ 5.1 Initialization phase
The construction of the starting kernel set $\Lambda^{0}$ and of the set of buckets $\mathcal{B}$ is carried out by Algorithm 3. The algorithm takes as input two positive parameters $K$ and $d$ (see Table 1) and returns both the sets $\Lambda^{0}$, and $\mathcal{B}$, as well as a starting solution $x^{0}$ together with its value $w^{0}.$
Input: Graph $\mathcal{G}=(V,E),$ conflict graph $\hat{\mathcal{G}}$, two scalars $K>0,d>0$.
Output: Initial kernel set $\Lambda^{0}$, initial bucket set $\mathcal{B}$, the first MSTC solution found $x^{0}$, and its value $w^{0}.$
1
2Compute the optimal solution $\bar{x}^{\prime}$ of the LP-MSTC relaxation.
3 Compute the optimal solution $\bar{x}^{\prime\prime}$ of the $\overline{\text{LP-MSTC}}$ relaxation.
4 if conflicts($(\{i:\bar{x}_{i}^{\prime}>0\})$ $<$ conflicts($(\{i:\bar{x}_{i}^{\prime\prime}>0\})$ then
5       $\bar{x}\leftarrow\bar{x}^{\prime}$
6else
7       $\bar{x}\leftarrow\bar{x}^{\prime\prime}$
8
[…]
10 Sort indices $i\in N$ in non-increasing order of $\bar{x}_{i}$ values.
11 $K\leftarrow\min\{K,|N|\}$.
12 $N_{K}\leftarrow N[1:K]$.
13 Compute the sets $\bar{E}$ and $S$ through the StartingSolution algorithm, with starting edge set $S^{0}:=N_{K}$.
14 Compute the independent set $I$ over the graph $\mathcal{H}:=\hat{\mathcal{G}}[S]$, through the IndependentSet algorithm.
15 $\Lambda^{0}\leftarrow I\cup\bar{E}$, and $L\leftarrow E\setminus\Lambda^{0}$.
16 Sort indices $i\in L$ according to Procedure 1.
17 if $|\Lambda^{0}|<K$ then
18      $\gamma=K-|\Lambda^{0}|$
19       $\bar{L}=L[1:\gamma]$
20       $\Lambda^{0}\leftarrow\Lambda^{0}\cup\bar{L}$
21       $L\leftarrow L\setminus\bar{L}$
22 $k\leftarrow 1$
23 while $L\neq\emptyset$  do
24       $d\leftarrow\min\{d,|L|\}$
25       $B_{k}\leftarrow L[1:d]$
26       $\mathcal{B}\leftarrow\mathcal{B}\cup B_{k}$
27       $L\leftarrow L\setminus B_{k}$
28       $k\leftarrow k+1$
29 Solve MSTC$(\Lambda^{0})$.
30 if MSTC$(\Lambda^{0})$ is feasible then
31      Let $x^{0}$ be the optimal solution of MSTC$(\Lambda^{0})$, and $w^{0}$ its optimal value.
32else
33      $x^{0}\leftarrow 0$, and $w^{0}\leftarrow+\infty$.
return $\Lambda^{0},\mathcal{B},x^{0},w^{0}.$
Algorithm 3 KS: Initialization Phase
First, the optimal solutions of both the LP-MSTC relaxation and the $\overline{\text{LP-MSTC}}$ relaxation are computed (lines 33). Then, the relaxed solution that results in the greatest number of compatible variables (i.e., those that are more likely to form an independent set in the conflict graph) is selected using the condition in lines 33. Specifically, the solution with fewer conflicts among variables with positive values is chosen and denoted as $\bar{x}$. Next, the set $N$ is constructed by including the indices of the variables with positive values in $\bar{x}$ (line 3), and $K$ is set in such a way that it is not greater than the cardinality of $N$ (line 3)
The subset of the first $K$ elements of $N$ is defined as $N_{K}$ (line 3). In line 3, the StartingSolution algorithm proposed in Algorithm 2 is executed with $N_{K}$ as starting edge set $S^{0}.$ The heuristic returns two edge sets: $\bar{E}$ and $S$. The set $\bar{E}$ is either a conflict-free spanning tree, or an empty set. The set $S$ is a superset of the set $N_{K}$, which is used to define the graph $\mathcal{H}$ needed as input to the IndependentSet algorithm. Indeed, in line 3, a feasible solution $I$ of the maximum independent set problem on the conflict subgraph $\hat{\mathcal{G}}[S]$ is computed through the greedy procedure presented in Algorithm 1, by setting $\mathcal{H}=\hat{\mathcal{G}}[S]$. This procedure guarantees that the returned set $I$ does not generate cycles in the starting graph $\mathcal{G}$.
Given the two edge sets $I$ and $\bar{E}$, coming from the IndependentSet algorithm and the StartingSolution algorithm, respectively, in line 3 the initial kernel set $\Lambda^{0}$ is populated with the indices of the variables associated with the edges in $I\cup\bar{E}$, and the remaining indices are used to define set $L$, which is used first to complete the set $\Lambda^{0}$ if its cardinality is smaller than $K$, and then to populate the buckets. The elements of set $L$ are sorted in line 3 according to the following procedure:
Procedure 1
The elements of $L$ are sorted sequentially according to the following criteria: $(i)$ non-decreasing order of the number of conflicts with the edges in $\Lambda^{0}$; $(ii)$ non-increasing order of value in the solution of LP-MSTC; $(iii)$ non-increasing order of the reduced cost coefficients. If two variables are tied according to a given criterion, the next criterion is used to determine priority. If there is still a tie after all three criteria are applied, the variable with the larger index $i$ is given priority.